OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x - 1 y = 4x - 5 y = 4x - 10 y = 4x + 1 What is the x-intercept and y-intercept of the graph of 3x - 4y = 12? x-intercept = (-4, 0); y-intercept = (0, 3) x-intercept = (4, 0); y-intercept = (0, -3) x-intercept = (0, -3); y-intercept = (4, 0) x-intercept = (0, 3); y-intercept = (-4, 0)

OpenStudy (anonymous):

If f(x) = x^2 + 2, find f(4). 10 18 2 6

OpenStudy (hba):

\[1)y-y_1=m(x-x_1)\] 2)For x-intercept plug in y=0 For y-intercept plug in x=0 3)Just plug in 4 in place of x